OpenStudy (anonymous):

How would I do 9+(-3) (negative 6) (please give steps)

OpenStudy (anonymous):

there is a law called PEMDAS.

OpenStudy (anonymous):

9-3 6

OpenStudy (anonymous):

i mean negative 6

OpenStudy (anonymous):

when u multiply + to - its a -

OpenStudy (amistre64):



OpenStudy (anonymous):

do everything in the parenthisis. which just holds a -3. then . then you just have 9-3.

OpenStudy (anonymous):

the -3 means negative three

OpenStudy (anonymous):

Hey @tic-tock - so what part are you having trouble with? Basically you have an equation where you are 'adding' a negative number to a positive number, right? Hence the: \[9+(-3)\]

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

So... given that, as others have shown, that is the same as simply subtracting. So, \[9+(-3) = 9-3\]

OpenStudy (anonymous):

negative 3 @cshalveny

OpenStudy (anonymous):

Sorry, negative 3 for what exactly?

OpenStudy (anonymous):

its negative 3

OpenStudy (anonymous):

9+negative 3

OpenStudy (anonymous):

Well, do I have your original equation, I *had* 9+ (-3) - right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and what I showed you was that 9 + a negative 3 is the same as 9 minus 3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so it is basically keep-change-change

OpenStudy (anonymous):

It only gets complicated when you subtract a negative, in those cases the sign changes. So, 9 - (-3) = 9+3

OpenStudy (anonymous):

so the answer would be 12

OpenStudy (anonymous):

No, in your problem remember that it equals 9-3, right?

OpenStudy (anonymous):

oh ya so the answer would be 6

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

Thanks so much cshalvey